OpenStudy (anonymous):

find the derivative of In(x/2)

OpenStudy (anonymous):

Well the derivative of ln(x) is 1/x So, that is half the battle right there. Then you must multiply by the chain rule. So you get (1/x)*(1/2)

OpenStudy (raden):

u are wrong, @cherio12 y=In(x/2) let u=x/2 ------> du/dx = 1/2 y=ln(u) ---------> dy/du = 1/u = 1/(x/2) = 2/x dy/dx = du/dx * dy/du dy/dx = 1/2 * 2/x = 1/x

OpenStudy (anonymous):

thanks

OpenStudy (raden):

yw

zepdrix (zepdrix):

Yah you had the right idea with the chain rule @cherio12 :) But you just forgot to place the ENTIRE argument of the natural log into the denominator, not just the x.\[\large \large \frac{d}{dx}\ln\left(\frac{x}{2}\right)=\frac{1}{\left(\frac{x}{2}\right)}\frac{d}{dx}\frac{x}{2}\] \[\large =\frac{2}{x}\cdot \frac{1}{2} \quad = \quad \frac{1}{x}\]